OpenStudy (anonymous):

NEED FAST HELP will give medal to first good answer Simplify 1. 2(8d+5) 2. 6y+4+8y 3. 6f+f 4. 6+8m-5m

OpenStudy (anonymous):

ANYBODY!!

OpenStudy (anonymous):

@mathmath333??? HELP

OpenStudy (mathmath333):

2(8d+5)=2*8d+2*5=16d+10

OpenStudy (mathmath333):

6y+4+8y=6y+8y+4 =y(6+8)+4 =12y+4

OpenStudy (mathmath333):

6f+f=f(6+1)=7f

OpenStudy (mathmath333):

6+8m-5m=6+m(8-5)=6+3m